$SPAD/src/input intdeq.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool intdeq.output
)set message test on
)set message auto off
)clear all
--S 1 of 7
y := operator y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 1
--S 2 of 7
deq := differentiate(y x, x, 2) + 2*w[0]*differentiate(y x, x) + w[0]**2*y x
--R
--R
--R         ,,          ,        2
--R   (2)  y  (x) + 2w y (x) + w  y(x)
--R                   0         0
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 7
sol:= solve(deq = sin (w*x), y, x=0,[0,0])
--R
--R
--R                                                      - w x
--R            2     2               3     2                0
--R        (- w  + w  )sin(w x) + ((w  + w  w)x + 2w w)%e      - 2w w cos(w x)
--R                 0                     0         0              0
--R   (3)  -------------------------------------------------------------------
--R                                  4      2 2     4
--R                                 w  + 2w  w  + w
--R                                        0       0
--R                                         Type: Union(Expression(Integer),...)
--E 3
--S 4 of 7
work:= sol *sin(w*x)

--R
--R
--R   (4)
--R           2     2         2
--R       (- w  + w  )sin(w x)
--R                0
--R     +
--R                               - w x
--R           3     2                0
--R       (((w  + w  w)x + 2w w)%e      - 2w w cos(w x))sin(w x)
--R                0         0              0
--R  /
--R      4      2 2     4
--R     w  + 2w  w  + w
--R            0       0
--R                                                    Type: Expression(Integer)
--E 4
--S 5 of 7
integrate(work,x)
--R
--R
--R   (5)
--R                                                - w x
--R                  4      3 2       4      2 2      0      4     4
--R         (((- 2w w  - 2w  w )x + 2w  - 6w  w )%e      + (w  - w  )cos(w x))
--R                0       0                0                     0
--R      *
--R         sin(w x)
--R     +
--R                                            - w x
--R             5      2 3         3              0         3      3          2
--R       ((- 2w  - 2w  w )x - 8w w )cos(w x)%e      + (2w w  + 2w  w)cos(w x)
--R                   0          0                        0       0
--R     +
--R           5     4
--R       (- w  + w  w)x
--R                0
--R  /
--R       7      2 5      4 3      6
--R     2w  + 6w  w  + 6w  w  + 2w  w
--R             0        0        0
--R                                         Type: Union(Expression(Integer),...)
--E 5
--S 6 of 7
D(%,x)-work
--R
--R
--R          2     2         2     2     2         2    2     2

--R        (w  - w  )sin(w x)  + (w  - w  )cos(w x)  - w  + w
--R               0                     0                    0
--R   (6)  ----------------------------------------------------
--R                           4      2 2      4
--R                         2w  + 4w  w  + 2w
--R                                 0        0
--R                                                    Type: Expression(Integer)
--E 6
--S 7 of 7
simplify %
--R
--R
--R   (7)  0
--R                                                    Type: Expression(Integer)
--E 7
)spool
)lisp (bye)

———-

References^{[1] nothing}

6